OpenStudy (anonymous):

Please help with limits question. Lim as x approaches 0 , (3/x -3/x^2-x) I know that you use conjugate and the answer is 3, but I can't figure out the steps to get to that answer. Please help!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is a bit hard to read

OpenStudy (misty1212):

\[\frac{3}{x}-\frac{3}{x^2-x}\]?

OpenStudy (misty1212):

no that can't be it if the answer is 3 ...

OpenStudy (anonymous):

Hi! Yes ma'am..oh the bottom is adding here, it said \[\lim_{x \rightarrow 0} \frac{ 3 }{ x } -\frac{ 3 }{ x^2+x }\]

OpenStudy (misty1212):

oh i see add

OpenStudy (misty1212):

or rather "subtract"

OpenStudy (misty1212):

you get after cancelling \[\frac{3}{x+1}\] it is all algebra pretty much no conjugates needed

OpenStudy (misty1212):

ma'am lol

OpenStudy (misty1212):

let me know if this is clear or not i think you got off on the wrong track when you tried multiplying by the conjugate it will give you a big unnecessary mess

OpenStudy (anonymous):

I know what you get after cancelling. I'm confused how you get there. lol